OpenStudy (anonymous):

Solve 3 log6 2 + 2 log6 3 – log6 x = 2

OpenStudy (mertsj):

\[\log_{6}8+\log_{6}9-\log_{6}x=2 \]

OpenStudy (mertsj):

\[\log_{6}\frac{72}{x} =2\]

OpenStudy (mertsj):

\[6^2=\frac{72}{x}\]

OpenStudy (mertsj):

You can finish it